A 2170 kg car traveling at 12.2 m/s collides with a 2940 kg car that is initally at rest at a stoplight. The cars stick together and move 2.02 m before friction causes them to stop. Determine the coefficient of kinetic friction between the cars and the road, assuming that the negative acceleration is constant and all wheels on both cars lock at the time of impact.
Answer #1

Using law of conservation of momentum,

m1 v1 + m2 v2 = (m1 + m2) v

2170 x 12.2 + 0 = (2170 + 2940) v

Velocity, v = 5.18 m/s

Using 3rd equation of motion,

v^2 = u^2 + 2as

5.18^2 = 0 + 2 x a x 2.02

Acceleration, a = 6.64 m/s^2

Frictional force, F = (2170 + 2940) x 6.64 = 33930.4 N

Coefficient of friction, u = 33930.4/(2170 + 2940) x 9.8

u = 0.678

Comment in case any doubt please rate my answer....